# Developing Safe and Responsible Large Language Models - A Comprehensive Framework

### v1

@@ 4.1 Training Details and Hyper-Parameters
The SR model is fine-tuned on a single A40 GPU with support from 4 CPU cores, employing PEFT and 4-bit quantization via QLORA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. Training was constrained to 1 epoch (with trials up to 5) to avoid over-fitting, similar to vanilla LLama2-7B experiences. We used a batch size of 16 for training, 8 for evaluation, saved checkpoints every 25 steps with an early stopping after 3, and set the learning rate to 2e-4 and utilized paged AdamW optimizer “paged_adamw_32bit”[dettmers_qlora_2023]. The max sequence length was limited to 1024 for faster inference, as well as greedy decoding strategy . The detailed hyper-parameters are given below:
@@
Carbon Footprint : To measure the environmental impact of training the SR model, the PEFT setup using one A40 GPU and four CPUs for 50 minutes had an energy use of 0.53 kWh and emitted 0.21 kgCO2e . This carbon footprint [dodge_measuring_2022] is notably low, especially when contrasted with more demanding tasks, such as a dense (full) fine-tuning, or training LLama2, which produced 539 tCO2eq, fully offset by Meta’s sustainability efforts. The calculations for carbon footprinting are given in C.
@@ 4.2 Evaluation Datasets
In-house Testsets: Our in-house test set, derived from our CMD dataset, consists of 6,000 entries, showing a wide array of content sensitivities and emotional tones. It includes instances classified by bias (1,732 biased, 4,268 unbiased), toxicity (3,612 non-toxic, 1,588 mildly toxic, 800 highly toxic), and sentiment (2,708 negative, 2,511 neutral, 780 positive), alongside a categorization based on harm potential (4,245 low, 1,180 medium, 574 high) and safety (3,108 unsafe, 2,892 benign). This diversity facilitates comprehensive analyses of safety interventions’ effectiveness across multiple content dimensions.
Additionally, we created a Counterfactual prompt-based test set, with 520 unique entries, is structured to probe biases related to gender, race, and religion through varied contextual modifications of base (neutral) sentences. This design enables a detailed examination of LLM biases and their capacity to handle complex, real-world scenarios involving critical social dimensions.
@@
Toxigen [hartvigsen_toxigen_2022]: Utilizing Toxigen v2 [hosseini_empirical_2023], refined to diminish noise by excluding sentences with annotator demographic disagreement, comprising 430 examples across 13 demographics.
@@
[…]
@@
Stereoset [nadeem_stereoset_2021]: Targets stereotype evaluation across demographics, with 8,498 samples spanning gender, race, profession, and religion.
@@
RedditBIAS [barikeri_redditbias_2021]: Analyzes bias within Reddit posts, with 1216 samples across gender, race, religion, and queerness.
@@
HolisticBIAS [smith_im_2022]: Features crowd-sourced prompts over 13 demographics, yielding 650 samples from 600 descriptors.
The details of these test sets are in D.
@@ 4.3 Baselines
Probability-based Baselines: These baselines rely on statistical probabilities and distributional information. They often serve as simple reference points for comparison. In our probability-based baseline approach, we use models like T5 [raffel_exploring_2019], Flan T5 [chung_scaling_2022], and BART [lewis_bart_2019] in various configurations to score the likelihood (probabilities) of content being unsafe versus benign.
Prompt-based Baselines: These baselines involve using specific prompts or instructions to guide model behavior. They can be tailored to address specific tasks or requirements. For LLMs, our study involves fine-tuned Llama2 variants, GPT-based models, and Falcon-instruct-7B in prompt-based settings with in-context (few-shot) learning (using safety prompts as in F.1).
The details of these baselines are given in E.
@@ 4.4 Evaluation Metrics
, and other categories from our ‘Safety risk taxonomy’ using our fine-tuned classifiers on each category to generate probability scores. Detailed information on these fine-tuned classifiers for bias, harm, and sentiment analysis is provided in E.1.
The Perspective API [perspective_api_perspective_2024] is employed for obtaining toxicity probabilities.
LLM Evaluation: Additionally, we use the GPTScore [fu_gptscore_2023] for evaluating LLMs and measuring the GPT model’s confidence in assessing unsafe text generations.
For content moderation, we utilize OpenAI’s moderation API [openai_moderation_2024] to obtain confidence scores. These scores, with a threshold of >0.5, are used to classify content as unsafe or benign.
Fairness Metrics: We adopt the fairness metrics from Steroeset [nadeem_stereoset_2021], which are:
@@
Content Diversity and Style Metrics: To measure the stylistic features in the texts post-safety interventions.
Partial Gen Bias metric [smith_im_2022] refers to a measure of how well a language model generates text that is partially biased or subtly reflects demographic imbalances. It assesses the model’s ability to produce content that exhibits bias without being overtly toxic or offensive.
CLEN (Content-Length Entropy Normalization) [miller_parlai_2017] is a metric employed to measure the diversity of sentence lengths in generated text, with a higher CLEN value signifies that the model contributes to stylistic variety. In the context of a style classifier, a higher CLEN value associated with a positive trait indicates that the sentence aligns more coherently with that style. For instance, in a benign variation, a higher CLEN value would imply greater stylistic consistency.
@@
T-tests are statistical tests that compare the means of two groups or a group mean against a standard to ascertain if observed differences are significant or due to chance [kim_t_2015].
We used One-Sample T-Test [ross_one-sample_2017] to evaluate the impact of safety interventions on texts, this test compares the mean characteristic (e.g., proportion classified as safe) before and after an intervention to a benchmark.
Further details on these metrics are given in E.2.
@@ 5 Results
5.1 Safety Evaluation
We evaluated SR using various state-of-the-art models on three test sets: our primary CMD test set, complemented by two prompt-based datasets: Toxigen and BOLD. Three classification setups were employed: a RoBERTa-based classifier fine-tuned for the ‘unsafe’ (ANNOT. ) label, the Perspective API (providing toxicity level probability scores), and the OpenAI moderation tool (detecting content violations and assigning confidence scores). We considered ‘unsafe’ label percentages, toxic generations based on probability scores (threshold > 0.5), and content violations from each classification setup. T5 and BART are used as fine-tuned, probability-score-based baselines; while Llama2, its variants, Falcon, and GPT-based models operate in an adapted few-shot setup. An example of this prompt with few-shot demonstration is given in F.1. Analysis measured ‘unsafe’ labels, ’toxic’ outputs, and ’violations’ as percentages (%) – drawing this setup inspiration from the original Llama2 work [touvron_llama_2023].
@@ 5.1 Safety Evaluation
The results presented in Table 5 highlights the effectiveness of SR in mitigating the generation of unsafe content, surpassing or equalling the performance of vanilla Llama2-chat models post additional safety fine-tuning. We observe that GPT-4 demonstrates superior content moderation capabilities compared to GPT-3.5, while the larger variants of Llama2 models exhibit a tendency towards marginal instances of unsafe content, consistent with prior observations in Llama2 and Llama-guard research. Despite careful fine-tuning efforts, T5 and BART models persistently exhibit relatively lower safety scores, indicating the persistent challenge of ensuring safety in generation tasks for these architectures.
@@ 5.2 Safety Evaluation Across Diverse Demographics
In this section, we show the safety evaluation of different models across different demographics. We used ToxiGen-RoBERTa [hartvigsen_toxigen_2022] to calculate the percentages of toxic text generation for each demographic group on the Toxigen data .
@@ 5.3 Safety Evaluation for Stereotypical Biases
We evaluate the performance of SR with different competitive models on StereoSet , targeting overall bias as well on across four domains. To delve into sentence-level biases, we opt for the Intrasentence task over the discourse-level Intersentence Test, aligning with our focus on detailed analysis of language generation and bias evaluation.
Table 7 shows SR shows best result in reducing biases and maintaining language quality, leading in SS and ICAT scores. GPT2-large exhibits notable stereotypical biases. SR best balances stereotype reduction and language performance, SR ICAT score are very close to Llama2-7b. Table 8 details bias results in gender, profession, race, and religion, highlighting race and religion as the most affected demographics. Nevertheless, SR efforts indicate progress in diminishing stereotypes, especially in gender and profession.
@@ 5.4 Safety Evaluation on Text Style Factors
Text style factors refer to the distinctive elements that influence the presentation and perception of written text, including vocabulary, sentence structure, tone, voice, and formatting [smith_controlling_2020]. These factors collectively define the unique character and readability of the text, impacting its engagement and effectiveness for the intended audience.
@@
To evaluate the effectiveness of safety measures implemented in SR, a controlled experiment was conducted, employing the ParlAI style classifier [smith_controlling_2020] to analyze the stylistic attributes of texts before and after safety interventions. This study aimed to determine whether these interventions led to significant changes in linguistic style and CLEN scores. The investigation was structured around a one-sample t-test, comparing the mean style scores of 16,602 textual instances against a hypothesized neutral value indicative of no unsafe generations. The null hypothesis () assumed no significant change in style due to the safety measures, whereas the alternative hypothesis () anticipated a discernible shift towards safer expressions.
@@ 5.4.1 Stylistic Variations Post-Safety Interventions
These two network diagrams illustrate contrasting collections of personality traits, each with a different focal point and emotional tone. The first network 3 (a) emphasizes more negative or challenging traits such as “Neurotic," “Hostile" and “Cruel," creating a sense of tension and difficulty within interpersonal dynamics. These terms interlink to form a complex web that suggests a personality model centered on conflict and struggle. In contrast, the second diagram 3(b) highlights positive and socially admirable qualities like “Caring," “Compassionate", and “Honest". This shows a sense of warmth and positive engagement post-safety intervention from SR , with traits that suggest nurturing, integrity, and enthusiasm. The central nodes in this network shows a character driven by empathy and understanding.
Table 9 showcases a comparison of Partial Generative Bias Scores before and after implementing safety interventions on texts, revealing a trend towards more affirmative styles post-intervention. In the pre-safety phase, the vanilla LLama2-7B model was utilized, while post-safety refers to SR, an iteration safety-tuned on the original Llama2-7B. The variability in Partial Generative Bias Scores serves as an indicator of bias levels, with the data demonstrating that styles exhibiting greater variance in Coherent Length Normalization (CLEN) possess higher bias scores. Within the framework of style classification, a higher CLEN value tied to a positive trait suggests that the text is more cohesively aligned with the said style. For example, in a benign variation, a higher CLEN value denotes a more consistent style.
@@
As evident in Table 9, prior to safety measures, the model predominantly generated text in a “Resentful" style, evidenced by its highest Partial Gen Bias Score of 1.000. Additional pre-safety styles showing notable bias included “Neurotic", “Meticulous" and “Outrageous", which could reflect various potentially problematic biases in text generation.
After safety interventions, there is a marked transition towards styles deemed more positive and constructive. “Scholarly" records the highest Partial Gen Bias Score of 0.758, with “Businesslike" reaching the peak score of 1.000, denoting an absolute shift in the model’s text generation preference. The styles “Articulate" and‘Eloquent" also receive high scores, indicating a model now predisposed to generating text that is more polished and well-articulated.
Overall, the analysis indicates that post-safety interventions, SR, exhibits a more profound alignment with constructive styles. This marks a considerable enhancement in the model’s generative bias, steering it towards generating content that is better suited for broader applications.
@@ 5.5 Human Evaluation
We assess SR variants for their ability to minimize harm, bias, toxicity, and stereotypes, targeting content neutrality and inclusivity through human evaluation. The model variations include Safe_PEFT-1_ep, fine-tuned for one epoch to examine minimal parameter adjustment effects on bias mitigation; Safe_PEFT-5_ep, which extends PEFT to five epochs to potentially enhance neutral content creation; Safe_Dense-IFT, subjected to comprehensive dense fine-tuning over five epochs for in-depth post-safety intervention; and Safe_Dense-IT, integrating dynamic task determination for adaptable bias mitigation strategies. These variants are tested against criteria of content neutrality, inclusivity, and respectful interaction, using a prompt designed to evaluate their output’s adherence to these standards. A team of five evaluators from varied backgrounds in ethics, social science, and computational linguistics conducts a rapid, blind assessment of 100 examples, chosen for their ability to robustly represent the performance of each model variant. The results are compiled in Table LABEL:tab:human, summarizing the human evaluation outcomes.
The evaluation of SR model variations in Table LABEL:tab:human shows diverse approaches to enhancing text neutrality and inclusivity. The Safe_PEFT-1_ep model performed best in removing unsafe generations (through example emphasizing diversity and skill in tech entrepreneurship), which led to the highest safety and language understanding scores. Safe_PEFT-5_ep brought Spanish phrases that was not needed, and it also slightly missed the mark on full inclusivity. Safe_Dense-IFT presented diversity in tech but concluded its response abruptly, affecting its evaluation scores. Safe_Dense-IT, despite an unnecessary apology for bias, broadened the demographic narrative effectively, scoring well in both safety and understanding. Overall, the results showcase progress in LLM generations to produce inclusive and respectful content, with Safe_PEFT-1_ep showing best for its comprehensive and inclusive approach.

### v2

@@ 3.1 Training Details and Hyper-Parameters
The SR model is fine-tuned on a single A40 GPU with support from 4 CPU cores, employing PEFT and 4-bit quantization via QLORA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. Training was constrained to 1 epoch (with trials up to 5) to avoid over-fitting, similar to vanilla LLama2-7B experiences. We used a batch size of 16 for training, 8 for evaluation, saved checkpoints every 25 steps with an early stopping after 3, and set the learning rate to 2e-4 and utilized paged AdamW optimizer “paged_adamw_32bit”Dettmers et al. (2023). The max sequence length was limited to 1024 for faster inference, as well as greedy decoding strategy . The detailed hyper-parameters are given below:
@@
Carbon Footprint : To measure the environmental impact of training the SR model, the PEFT setup using one A40 GPU and four CPUs for 50 minutes had an energy use of 0.53 kWh and emitted 0.21 kgCO2e . This carbon footprint Dodge et al. (2022) is notably low, especially when contrasted with more demanding tasks, such as a dense (full) fine-tuning, or training LLama2, which produced 539 tCO2eq, fully offset by Meta’s sustainability efforts. The calculations for carbon footprinting are given in D.
@@ 3.2 Evaluation Datasets
In-house Testsets: Our in-house test set is derived from our CMD dataset, consists of 6,000 entries. It includes instances classified by bias (1,732 biased, 4,268 unbiased), toxicity (3,612 non-toxic, 1,588 mildly toxic, 800 highly toxic), and sentiment (2,708 negative, 2,511 neutral, 780 positive), alongside a categorization based on harm potential (4,245 low, 1,180 medium, 574 high) and safety (3,108 unsafe, 2,892 benign). This diversity facilitates comprehensive analyses of safety interventions’ effectiveness across multiple content dimensions.
Additionally, we present a Counterfactual prompt-based test set, with 520 unique entries, is structured to probe biases related to gender, race, and religion through varied contextual modifications of base (neutral) sentences.
@@
BOLD Dhamala et al. (2021): A Wikipedia-based, prompt-driven dataset spanning four demographic groups (race, gender, religion, and profession) with 7200 samples.
@@
Stereoset Nadeem et al. (2021): Targets stereotype evaluation across demographics, with 8,498 samples spanning gender, race, profession, and religion.
@@
RedditBIAS Barikeri et al. (2021): Analyzes bias within Reddit posts, with 1216 samples across gender, race, religion, and queerness.
@@
HolisticBIAS Smith et al. (2022): Features crowd-sourced prompts over 13 demographics, yielding 650 samples from 600 descriptors.
@@ 3.3 Baselines
Probability-based Baselines: These baselines rely on statistical probabilities and distributional information. In our probability-based baseline approach, we use models like T5 Raffel et al. (2019), Flan T5 Chung et al. (2022), and BART Lewis et al. (2019) in various configurations to score the likelihood (probabilities) of content being unsafe versus benign.
Prompt-based Baselines: These baselines involve using specific prompts or instructions to guide model behavior. They can be tailored to address specific tasks or requirements. For LLMs, our study involves fine-tuned Llama2 variants, GPT-based models, and Falcon-instruct-7B in prompt-based settings with in-context (few-shot) learning (using safety prompts as in E.3).
The T5 and its advanced variant FLan-T5 were employed in configurations up to 770 million parameters, fine-tuned with learning rates from to , batch sizes, and up to 20 epochs. The BART-large model, leveraging bidirectional and autoregressive capabilities, followed similar fine-tuning specifics aimed at task-specific optimization. Llama 2-7B an 13B from Meta AI and the Generative Pre-trained Transformer (GPT) series, including GPT2-large, DialoGPT, GPT-3.5 Turbo, and GPT-4, provided a range from 774 million to hundreds of billions of parameters for deep language generation and understanding explorations. Lastly, Falcon-7B-Instruct, a model fine-tuned on chat and instruction datasets, emphasized safety and robust performance.
@@ 3.4 Evaluation Metrics
, and other categories from our ‘Safety risk taxonomy’ using our fine-tuned classifiers on each category to generate probability scores. Detailed information on these fine-tuned classifiers for bias, harm, and sentiment analysis is provided in Appendix E.
The Perspective API API (2024) is employed for obtaining toxicity probabilities.
LLM Evaluation: Additionally, we use the GPTScore Fu et al. (2023) for evaluating LLMs and measuring the GPT model’s confidence in assessing unsafe text generations.
For content moderation, we utilize OpenAI’s moderation API OpenAI (2024) to obtain confidence scores. These scores, with a threshold of >0.5, are used to classify content as unsafe or benign.
Fairness Metrics: We adopt the fairness metrics from Steroeset Nadeem et al. (2021), which are:
@@
Content Diversity and Style Metrics: To measure the stylistic features in the texts post-safety interventions (in benign generations from LLM), we use:
Partial Gen Bias metric Smith et al. (2022) that refers to a measure of how well a language model generates text that is partially biased or subtly reflects demographic imbalances. It assesses the model ability to produce content that exhibits bias without being overtly toxic or offensive.
CLEN (Content-Length Entropy Normalization) Miller et al. (2017) is a metric employed to measure the diversity of sentence lengths in generated text, with a higher CLEN value signifies that the model contributes to stylistic variety. In the context of a style classifier, a higher CLEN value associated with a positive trait indicates that the sentence aligns more coherently with that style. For instance, in a benign variation, a higher CLEN value would imply greater stylistic consistency.
@@
T-tests are statistical tests that compare the means of two groups or a group mean against a standard to ascertain if observed differences are significant or due to chance Kim (2015).
We used One-Sample T-Test Ross and Willson (2017) to evaluate the impact of safety interventions on texts, this test compares the mean characteristic (e.g., proportion classified as safe) before and after an intervention to a benchmark.
Further details on these metrics are given in Appendix E.
@@ 4 Results
4.1 Overall Safety Evaluation
We evaluated SR using various state-of-the-art language models on three test sets: our primary CMD test set, complemented by two prompt-based datasets: Toxigen and BOLD. Three classification setups were employed: our fine-tuned RoBERTa-based classifier for the ‘unsafe’ (ANNOT. ) label, the Perspective API (providing toxicity level probability scores), and the OpenAI moderation tool (detecting content violations and assigning confidence scores).
We considered ‘unsafe’ label percentages, toxic generations based on probability scores (threshold > 0.5), and content violations from each classification setup. T5 and BART are used as fine-tuned, probability-score-based baselines; while Llama2, its variants, Falcon, and GPT-based models operate in an adapted few-shot setup. An example of this prompt with few-shot demonstration is given in E.3. Analysis measured ‘unsafe’ labels, ’toxic’ outputs, and ’violations’ as percentages (%) – drawing this setup inspiration from the original Llama2 work Touvron et al. (2023).
@@ 4.1 Overall Safety Evaluation
The results presented in Table 4 highlights the effectiveness of SR in mitigating the generation of unsafe content, performing better or equal to the performance of vanilla Llama2-chat models post additional safety fine-tuning. We observe that GPT-4 demonstrates superior content moderation capabilities compared to GPT-3.5, while the larger variants of Llama2 models exhibit a tendency towards marginal instances of unsafe content, consistent with prior observations in Llama2 and Llama-guard research. Despite careful fine-tuning efforts, T5 and BART models persistently exhibit relatively lower safety scores, indicating the persistent challenge of ensuring safety in generation tasks for these architectures.
@@ 4.2 Safety Evaluation Across Diverse Demographics
In this section, we show the safety evaluation of different models across different demographics. We used ToxiGen-RoBERTa Hartvigsen et al. (2022) to calculate the percentages of toxic text generation for each demographic group on the Toxigen data .
@@ 4.3 Safety Evaluation for Stereotypical Biases
We evaluate the performance of SR with different competitive models on StereoSet , targeting overall bias as well on across four domains. To delve into sentence-level biases, we opt for the Intrasentence task over the discourse-level Intersentence Test, aligning with our focus on detailed analysis of language generation and bias evaluation. Table 6 shows SR shows best result in reducing biases and maintaining language quality, leading in SS and ICAT scores. GPT2-large exhibits notable stereotypical biases. SR best balances stereotype reduction and language performance, SR ICAT score are very close to Llama2-7b.
@@
Table 7 details bias results in gender, profession, race, and religion, highlighting race and religion as the most affected demographics.
@@ 4.4 Safety Evaluation on Text Style Factors
Text style factors refer to the distinctive elements that influence the presentation and perception of written text, including vocabulary, sentence structure, tone, voice, and formatting Smith et al. (2020). These factors collectively define the unique character and readability of the text, impacting its engagement and effectiveness for the intended audience.
@@
To evaluate the effectiveness of safety measures implemented in SR, a controlled experiment was conducted, employing the ParlAI style classifier Smith et al. (2020) to analyze the stylistic attributes of texts before and after safety interventions. This study aimed to determine whether these interventions led to significant changes in linguistic style and CLEN scores. The investigation was structured around a one-sample t-test, comparing the mean style scores of 16,602 textual instances against a hypothesized neutral value indicative of no unsafe generations. The null hypothesis () assumed no significant change in style due to the safety measures, whereas the alternative hypothesis () anticipated a discernible shift towards safer expressions.
@@ 4.4.1 Stylistic Variations Post-Safety Interventions
Figure 4 illustrate contrasting collections of personality traits, each with a different focal point and emotional tone. The first network 4 (a) emphasizes more negative or challenging traits such as “Neurotic," “Hostile" and “Cruel," creating a sense of tension and difficulty within interpersonal dynamics. These terms interlink to form a complex web that suggests a personality model centered on conflict and struggle. In contrast, the second diagram 4 (b) highlights positive and socially admirable qualities like “Caring," “Compassionate", and “Honest". This shows a sense of warmth and positive engagement post-safety intervention from SR , with traits that suggest nurturing, integrity, and enthusiasm. The central nodes in this network shows a character driven by empathy and understanding.
Table 8 presents a comparison of Partial Generative Bias Scores before and after safety interventions, showing a shift towards more positive styles after intervention. The table focuses on key styles for conciseness. Initially, the Llama2-7B model was used, and after safety measures, the SR version, a safety-enhanced Llama2-7B, was applied. A higher CLEN value indicates stronger alignment with the positive trait being measured. For instance, a high CLEN value associated with a positive trait like ’scholarly’ indicates a more consistently positive style.
@@
Results in Table 8 prior to our safety measures shows that the model predominantly generated text in many negative traints like “Resentful", "Outrageous", and such styles. After safety interventions, there is a transition towards styles deemed more positive and constructive. For example, “Businesslike", “Scholarly" ,“Articulate" and‘Eloquent" receive high scores, indicating a model now predisposed to generating text that is more polished and well-articulated. Overall, the analysis indicates that post-safety interventions, SR, exhibits a more profound alignment with constructive styles.
@@ 4.5 Human Evaluation
We assess SR variants for their ability to minimize harm, bias, toxicity, and stereotypes, targeting content neutrality and inclusivity through human evaluation. The model variations include:
Safe_PEFT-1_ep, fine-tuned for one epoch to examine minimal parameter adjustment effects on bias mitigation.
Safe_PEFT-5_ep, which extends PEFT to five epochs to potentially enhance neutral content creation
Safe_Dense-IFT, subjected to comprehensive dense fine-tuning over five epochs for in-depth post-safety intervention.
Safe_Dense-IT, integrating dynamic task determination for adaptable bias mitigation strategies. These variants are tested against criteria of content neutrality, inclusivity, and respectful interaction, using a prompt designed to evaluate their output’s adherence to these standards.
A team of five evaluators from varied backgrounds in ethics, social science, and computational linguistics conducts a rapid, blind assessment of 100 examples, chosen for their ability to robustly represent the performance of each model variant. Main results of human evaluation are compiled in Table LABEL:tab:human, summarizing the human evaluation outcomes.
@@
The evaluation of SR model variations in Table LABEL:tab:human shows diverse approaches to enhancing text neutrality and inclusivity. The Safe_PEFT-1_ep model performed best in removing unsafe generations (through example emphasizing diversity and skill in tech entrepreneurship), which led to the highest safety and language understanding scores. Safe_PEFT-5_ep brought Spanish phrases that was not needed, and it also slightly missed the mark on full inclusivity. Safe_Dense-IFT presented diversity in tech but concluded its response abruptly, affecting its evaluation scores. Safe_Dense-IT, despite an unnecessary apology for bias, broadened the demographic narrative effectively, scoring well in both safety and understanding. Overall, the results showcase progress in LLM generations to produce inclusive and respectful content, with Safe_PEFT-1_ep (which is out default setting) showing best for its comprehensive and inclusive approach.